OpenStudy (anonymous):

1) WILL GIVE MEDAL AND FAN!!!! x2 – 4x + 4 a. (x + 4)2 b. (x + 2)2 c. (x – 4)2 d. (x – 2)2

rishavraj (rishavraj):

\[(a - b)^2 = a^2 + b^2 - 2ab\]

OpenStudy (anonymous):

so do you know what answer that would be? @rishavraj

rishavraj (rishavraj):

ooo yeah its D (x - 2)^2 = x^2 + 4 - 4x

OpenStudy (anonymous):

thanks @rishavraj

rishavraj (rishavraj):

yw :))))